3/7(14x-21)=1/2(6x+40) Please Answer as quick as you can. I get some but not all.
Answer by tutorcecilia(2152)

3/7(14x-21)=1/2(6x+40)
(3/7)(14x/1)-(3/7)(21)= (1/2)(6x/1)+ (1/2)(40/1) [Distribute]
(3)(2x)-(3)(3)= (3x)+ (20) [Cancel whereever possible]
6x-9= 3x+ 20 [Simplify]
6x-9= 3x+ 20 [Solve for x]
6x-3x-9= 3x-3x+ 20
3x-9+9=20+9
3x=29
3x/3=29/3
x=29/3 [Final answer]